# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ 4.4 Halo random walk and halo mass function
Figure 7 plots different dimensionless mass functions (log(f($\nu$)) (Appendix A.6)) compared to the mass function obtained from the Virgo simulation (solid blue). The PS mass function overestimates the mass in small haloes and underestimates the mass in large haloes. The JK mass function matches the simulation for large-mass haloes with a large deviation for small-mass haloes. The new double-$\lambda$ mass function (Eq. (35)) with the best fit of $\eta_{0}=0.76$ and $q=0.556$ matches both the simulation and the ST mass function for the entire range. With $\lambda_{2}={2/3}$ for the deposition range and $q=0.556$, we estimated $\lambda_{1}=0.815$ for the propagation range.
4.5 Formulating the energy cascade in mass space
@@
This section presents a quantitative description of the energy cascade in the halo mass space. To better describe the energy cascade, we first decompose the kinetic energy of the halo particles into two parts of different nature. In N-body simulations, every halo particle can be characterized by a mass $m_{p}$ and a velocity vector $\boldsymbol{\mathrm{v}}_{\boldsymbol{\mathrm{p}}}$. The particle velocity $\boldsymbol{\mathrm{v}}_{p}$ can be decomposed as (Xu, 2023a)
@@
namely, the halo mean velocity, $\boldsymbol{\mathrm{v}}_{h}=\langle\boldsymbol{\mathrm{v}}_{p}\rangle_{h}$ , and the velocity fluctuation, $\boldsymbol{\mathrm{v}}_{p}^{{}^{\prime}}$. Here, $\left\langle\right\rangle_{h}$ represents the average of all particles in the same halo, and $\boldsymbol{\mathrm{v}}_{h}$ represents the velocity of that halo. Consequently, the total kinetic energy $K_{p}$ of a given halo particle can be divided into $K_{p}=K_{ph}+K_{pv}$. Here, $K_{ph}=\boldsymbol{\mathrm{v}}_{h}^{2}/2$ (halo kinetic energy) is the contribution from the motion of entire haloes $\boldsymbol{\mathrm{v}}_{h}$ due to the inter-halo interaction of that particle with all other particles in different haloes and all out-of-halo particles. This part of the kinetic energy is related to interactions on large scales that are still in the linear regime.
The other part, $K_{pv}={\boldsymbol{\mathrm{v}}_{p}^{\prime}}^{2}/2$ (virial kinetic energy), is the contribution of the velocity fluctuation $\boldsymbol{\mathrm{v}}_{p}^{\prime}$ due to the intra-halo interaction of that particle with all other particles in the same halo. This part of the kinetic energy is from halo virialization and is due to interactions on a shorter distance and smaller scales in the non-linear regime. Similarly to the energy cascade associated with nonlinear interactions in turbulence, the energy cascade in dark matter flow focused on the cascade of the virial kinetic energy $K_{pv}$ due to the nonlinear interactions on small scales.
Similar to the cumulative mass function $\Lambda_{m}$ in Eq. (17), the cumulative kinetic energies ($\Lambda_{ph}$ and $\Lambda_{pv}$) represent the total kinetic energies $K_{ph}$ and $K_{pv}$ in all haloes greater than $m_{h}$, such that
[…]
@@
where $\overline{K_{pv}}$ is defined in Eq. (38), i.e. the specific virial kinetic energy in all haloes greater than $m_{h}$. Similar to Eq. (16), Eq. (39) describes the rate of transfer of specific virial kinetic energy ($K_{pv}$) from haloes below the scale $m_{h}$ to haloes above the scale $m_{h}$ at a rate of $\Pi_{pv}$.
Figure 8: The variation of the rate of energy cascade $\Pi_{pv}(m_{h},a)$ (Eq. (39)) with halo mass $m_{h}$ at different redshifts $z$ from Illustris-1-Dark simulation. A scale-independent constant rate of $\varepsilon_{u}$ can be identified in the propagation range for an inverse cascade of virial kinetic energy ${K_{pv}}$. That rate is also relatively independent of time and is around $\varepsilon_{u}=-2.5\times 10^{-7}m^{2}/s^{3}$. According to the virial theorem, there also exists a simultaneous direct cascade of potential energy from large to small scales.

### v4

@@ 3.1 Halo random walk and mass function
Figure 5 plots different analytical dimensionless mass functions (log(f($\nu$)) (Appendix C.6) compared to the mass function obtained from the Virgo simulation (solid blue). The PS mass function overestimates the mass in small haloes and underestimates the mass in large haloes. The JK mass function matches the simulation for large-mass haloes with a large deviation for small-mass haloes. The new double-$\lambda$ mass function (Eq. (28)) with the best fit of $\eta_{0}=0.76$ and $q=0.556$ matches both the simulation and the ST mass function for the entire range. With $\lambda_{2}={2/3}$ in the universal transition range and $q=0.556$, we estimate $\lambda_{1}=0.815$, an effective $\lambda$ that reflects the average waiting time for a given CDM spectrum.
3.2 Universal scaling laws in transition region
The mass cascade in the universal range dominant in gravity enables universal scaling laws. We will first quantify the parameter $\tau_{g}$ in that universal range. By definition, $\tau_{g}$ quantifies the mass accretion rate,
$\begin{split}&\frac{dm_{h}}{dt}=\frac{m_{p}}{\tau_{g}},\end{split}$ (29)
where $m_{p}$ denotes the characteristic mass increase per accretion event during smooth accretion (e.g., the mass associated with a single merger or a dark matter particle). This relation implies that, on average, the halo gains a mass $m_{p}$ over a time scale $\tau_{g}$. In practice, major mergers between haloes of comparable mass are infrequent, particularly during the early matter-dominated epoch when the halo number density is low. Equation (29) should therefore be interpreted as a coarse-grained effective accretion rate that captures smooth growth and the cumulative contribution of minor mergers and continuous inflow, rather than the accretion of major mergers. The waiting time $\tau_{g}\propto m_{h}^{-\lambda}$ depends on the halo mass through the parameter $\lambda$. Intuitively, the waiting time depends on the initial distribution of matter. The more negative $n$ leads to a higher large-scale power and more massive haloes, a lower small-scale power, less frequent merging, thus a large $\lambda$ and a longer waiting time $\tau_{g}$ for low-mass haloes. For a general CDM power spectrum, $\lambda$ varies with scale. However, a unique value of $\lambda$ can be derived for haloes around $m_{h}^{*}$ in the universal transition range.
In the transition range, since gravitational dynamics is dominant, the waiting time is inversely proportional to the halo potential $\tau_{g}\propto\Phi_{h}^{-1}$ (Xu:2023-Dark-matter-halo-mass-functions-and). This is also explained by the random walk of particles in haloes in Section 4.3 (Eq. (75)). Therefore, we can write the following for the haloes of $m_{h}\sim m_{h}^{*}$
$\begin{split}&\tau_{g}\propto\Phi_{h}^{-1}\propto(Gm_{h}/r_{h})^{-1},\\ &\rho_{h}=\frac{m_{h}}{\frac{4}{3}\pi{r_{h}}^{3}}=\Delta_{c}\rho_{DM}\propto a^{-3},\end{split}$ (30)
where $r_{h}$ is the size of the halo. The second equation relates the average halo density to the mean density $\rho_{DM}$, where $\Delta_{c}=18\pi^{2}$ is the critical density ratio. Solving these equations leads to:
A unique value of $\lambda=2/3$ can be predicted for haloes around $m_{h}^{*}$ in the universal transition range. Using the relation between $\lambda$ and $n$ (Eq. (25)), $\lambda=2/3$ corresponds to a universal spectrum index n=1n=-1 for the gravitational dominant transition range, as shown in Fig. 1.
Substituting $\tau_{g}\propto m_{h}^{-2/3}a$ into Eq. (29), with $\lambda=2/3$, the scaling relations for haloes with characteristic mass $m_{h}^{*}$ read
$r_{h}^{*}\propto t,\quad m_{h}^{*}\propto t\quad\textrm{and}\quad\tau_{g}^{*}\propto(m_{h}^{*})^{-2/3}a\propto t^{0}.$ (32)
Therefore, in the matter-dominant era, characteristic haloes accrete mass at a constant rate with a constant waiting time $\tau_{g}^{*}$ such that both mass $m_{h}^{*}$ and size $r_{h}^{*}$ are $\propto t\propto a^{3/2}$. Since the characteristic mass $m_{h}^{*}\propto a^{3/2}$, the comoving size $r_{hc}^{*}\propto a^{1/2}$ leads to an expanding universal range with an estimate of $\alpha_{x}=1/2$ in Eq. (3). Using solution (21), in the universal transition range, the halo mass function $f_{M}$ follows a simple power law (with $\lambda=2/3$ and $m_{h}^{*}\propto a^{3/2}$),
$\begin{split}&f_{M}(m_{h},z)\propto{m_{h}^{*}}^{\lambda-1}{m_{h}}^{-\lambda}\propto m_{h}^{-2/3}a^{-1/2}\end{split}$ (33)
Figure 6: The evolution of the characteristic halo mass $m_{h}^{*}(t)$ (red dashed) and total halo mass $M_{h}(t)$ (blue dashed) in the matter era, both can be numerically determined from N-body simulations (Xu:2023-Dark-matter-halo-mass-functions-and). Results from Illustris and Virgo simulations are presented as symbols, which are in agreement with the prediction of Eqs. (32) and (40), i.e. $m_{h}^{*}\propto a^{3/2}\propto t$ and $M_{h}\propto a^{1/2}\propto t^{1/3}$. At high redshift, before the universal spectrum being fully developed, $m_{h}^{*}\propto a^{3}$ follows the mass scale $M^{*}$ at $k_{NL1}$.
In this section, we identify the waiting time $\tau_{g}\propto m_{h}^{-2/3}a$ with $\lambda=2/3$ in the gravitationally dominant universal range, confirm the universal power spectrum nn=-1, and derive the scaling relations for the evolution of $m_{h}^{*}$. Figure 6 presents the time variation of $m_{h}^{*}$ from the Illustris and Virgo simulations. After the system reaches a statistically stable state with a fully developed universal spectrum, the characteristic mass is reached $m_{h}^{*}\propto a^{3/2}$. To derive the evolution of total mass $M_{h}(t)$ in all haloes, i.e., $M_{h}\propto a^{1/2}$, we need to discuss the scale-independent rate $\varepsilon_{m}$ of the mass cascade in the next section.
3.3 The scale-to-scale mass cascade in mass space
In this section, we present a quantitative description of the scale-to-scale mass transfer in the halo mass space, i.e. a mass cascade process. As shown in Fig. 4, the halo random walk in mass space generates mass and energy flux across scales. On average, it takes a waiting time $\tau_{g}(m_{h},z)$ for a given halo to merge and go to the next scale. For a group of $n_{h}$ haloes of the same mass $m_{h}$, the average waiting time for a merge event of any halo in that group can be found as $\tau_{h}(m_{h},z)=\tau_{g}/n_{h}$. A simple explanation is provided here.
Let us randomly pick one halo in that group to merge in every round and repeat until the given halo is picked. The probability of picking that given halo in the first round is $1/n_{h}$, and the corresponding waiting time for that specific halo is $\tau_{h}$. The probability of picking that given halo in the second round is $(n_{h}-1)/n_{h}*1/n_{h}$ (product of the probability not selected in the first round and the probability selected in the second round). The corresponding waiting time is $2\tau_{h}$. Therefore, the waiting time $\tau_{g}$ for a given halo is
$\begin{split}\tau_{g}=\sum_{k=1}^{\infty}\frac{k\tau_{h}}{n_{h}}\left(1-\frac{1}{n_{h}}\right)^{k-1}=\frac{\tau_{h}}{n_{h}}+...=n_{h}\tau_{h}.\end{split}$ (34)
Based on this description, the rate of mass cascade (or mass flux) in halo mass space can be conveniently defined as
$\begin{split}&\Pi_{m}(m_{h},z)=-\frac{m_{h}}{\tau_{h}}=-\frac{n_{h}m_{h}}{\tau_{g}}=-\frac{m_{g}}{\tau_{g}},\\ &\varepsilon_{m}(z)=\Pi_{m}(m_{h},z)\quad\textrm{for}\quad m_{h}\leq m_{h}^{*}.\end{split}$ (35)
The mass $m_{h}$ of the halo is transferred to a larger mass scale during a period of $\tau_{h}$. Here, $m_{g}=n_{h}m_{h}$ is the total mass for all $n_{h}$ haloes of the same mass $m_{h}$ (or the mass of the halo group), such that $\tau_{g}$ is the time required to transfer the mass of the entire group $m_{g}$ to a larger mass scale. The negative sign represents the inverse cascade from small to large scales, in contrast to the direct cascade.
Here, we introduce an important hypothesis implied by Eq. (35): the rate of the mass cascade is scale independent on scales $m_{h}<m_{h}^{*}$; that is, $\varepsilon_{m}$ is independent of the mass scale $m_{h}$ in that range. The mass and energy cascade drives the nonequilibrium system to establish a statistically steady state to continuously release the energy of the system and maximize entropy (Xu:2023-Maximum-entropy-distributions-of-dark-matter). When the dark matter flow reaches that statistically steady state, the rates of mass and energy cascade must be scale independent. If this is not the case, there would be a net accumulation of mass and energy on some intermediate-mass scale below $m_{h}^{*}$.
To validate this hypothesis by N-body simulations, we need to calculate the scale-to-scale transfer in simulations. We first introduce the mass flux ($\Pi_{m}$) across haloes of different sizes
$\begin{split}\Pi_{m}\left(m_{h},t\right)&=-\int_{m_{h}}^{\infty}\frac{\partial}{\partial t}\left[M_{h}\left(t\right)f_{M}\left(m,m_{h}^{*}\right)\right]dm,\\ &=-\frac{\partial}{\partial t}\left[\int_{m_{h}}^{\infty}M_{h}\left(t\right)f_{M}\left(m,m_{h}^{*}\right)dm\right]=-\frac{\partial\Lambda_{m}}{\partial t},\end{split}$ (36)
$\Lambda_{m}(m_{h},t)=\int_{m_{h}}^{\infty}M_{h}\left(t\right)f_{M}\left(m,m_{h}^{*}\right)dm.$ (37)
The total mass in all haloes can be obtained by setting $m_{h}\rightarrow 0$ in Eq. (37), i.e., $M_{h}(t)=\Lambda_{m}(m_{h}\rightarrow 0,t)$. The time derivative of $\Lambda_{m}$ describes the mass flux from all haloes below the scale $m_{h}$ to all haloes above the scale $m_{h}$, i.e., the rate of mass cascade $\Pi_{m}$ across the scale $m_{h}$.
Equation (36) describes the total mass flux induced by the binary merging involving three haloes (two merging haloes and one remnant halo), with at least one merging halo being below $m_{h}$ and the remnant halo above $m_{h}$. Under the assumption of smooth mass accretion, the dominant contribution to this total mass flux is the merging of $m$ and $m_{h}$ to produce a halo of mass $m+m_{h}$, such that Eq. (36) is equivalent to Eq. (35), especially in the early universe when halo number density is low. Appendix C.1 provides more discussion on this locality feature for halo interactions in mass space.
[…]
Figure 8: The variation of mass flux $\Pi_{m}\left(m_{h},z\right)$ from the Virgo simulation using halo group mass $m_{g}\left(a\right)$ at two different redshifts z (Eq. (36)). A scale-independent mass flux $\varepsilon_{m}\left(a\right)$ can be found for haloes smaller than the characteristic mass $m_{h}<m_{h}^{*}$. The negative mass flux $\varepsilon_{m}(z)<0$ indicates an inverse mass cascade from small to large scales. A mass propagation range with a scale-independent $\varepsilon_{m}$ is formed at around $z=0.3$, much later than the Illustris simulation in Fig. 7 due to a lower mass resolution.
Figure 8 presents the evolution of the mass flux function $\Pi_{m}\left(m,z\right)$ from the Virgo simulation. Similarly, a scale-independent mass flux $\varepsilon_{m}(z)$ can be clearly identified for the halo mass $m_{h}<m_{h}^{*}$. The propagation range with scale-independent mass flux $\varepsilon_{m}$ is formed around $z=0.3$ and gradually expands in mass space. The time to reach the statistical state with a scale-independent rate of cascade is dependent on the simulation resolution. The higher mass resolution in the Illustris simulation leads to an earlier formation of the propagation range. The simulation with a higher resolution establishes the statistically steady state at a higher redshift. Both simulations confirm a scale-independent rate of mass cascade.
An important and direct consequence of a scale-independent rate of the cascade $\varepsilon_{m}(a)$ is the total mass of a halo group $m_{g}$ invariant in time. Taking the derivative of Eq. (36) leads to
$\begin{split}&\frac{\partial\Pi_{m}}{\partial m_{h}}=-\frac{\partial}{\partial t}\left(\frac{\partial\Lambda_{m}}{\partial m_{h}}\right)=-\frac{\partial m_{g}}{\partial t}=0,\\ &m_{g}(m_{h})=M_{h}(t)f_{M}(m_{h},z)m_{p}=n_{h}m_{h}.\end{split}$ (38)
Here, $n_{h}$ is the number of haloes with the same mass $m_{h}$. The mass of the entire halo group $m_{g}$ is independent of the redshift $z$ due to the scale-independent mass cascade. That is, the mass flux into a halo group is equal to the mass flux out of the same group, so the mass of the group $m_{g}$ is a function of $m_{h}$ only and is constant over time. This is the so-called small-scale permanence that is independent of the cosmological model and initial conditions. As long as a statistical steady state is established in dark matter flow, the halo group mass $m_{g}$ at different redshifts $z$ always collapses into a redshift-independent $m_{g}(m_{h})$. Figure 9 demonstrates this concept from Illustris simulations. The mass of the halo group $m_{g}$ at different $z$ collapses into a simple power law $m_{g}(m_{h})\propto m_{h}^{-\lambda}$, with an effective $\lambda\approx$1.9 determined by the initial power spectrum of Illustris.
In addition, with mass injected from the smallest scale $m_{p}$ at a scale-independent rate of $\varepsilon_{m}$, the rate of the mass cascade also equals the rate of change in total mass $M_{h}(t)$ in all haloes. From the definitions of the rate of mass cascade $\varepsilon_{m}$ and the halo group mass $m_{g}$ (Eq. (35)), we have the following.
[…]
Combining Eq. (39) with Eq. (31) for $\tau_{g}\propto m_{h}^{-2/3}a$, additional scaling laws can be identified when the system reaches the statistically steady state with a scale-independent rate $\varepsilon_{m}$. In the universal transition range around mass $m_{h}^{*}$, when small-scale permanence is satisfied (time-independent $m_{g}$),
The rate of mass cascade $\varepsilon_{m}$ decreases with time as $\propto a^{-1}$, while the total halo mass $M_{h}(t)$ increases with time as $a^{1/2}$ (shown in Fig. 6). More generally, the mass of the halo group $m_{g}(m_{h})$ is independent of the redshift regardless of the initial conditions (small-scale permanence). In particular, the group mass $m_{g}^{*}$ and the total number of haloes $n_{h}^{*}$ for the characteristic haloes of mass $m_{h}^{*}$ are read (using Eqs. (38) and (33))
$\begin{split}&m_{g}^{*}\propto M_{h}(t)(m_{h}^{*})^{-1}m_{p}\propto a^{-1},\\ &n_{h}^{*}\propto{M_{h}(t)m_{p}}/{{m_{h}^{*}}^{2}}\propto a^{-5/2}.\end{split}$ (41)
Therefore, the total halo mass $M_{h}$ (or the total mass of dark matter) is known as long as we know the particle mass $m_{p}$ and the number $n_{h}^{*}$ for characteristic haloes. If $n_{h}^{*}$=1 at zz=0, $M_{h}\approx{m_{h}^{*}}^{2}/m_{p}$.
Figure 9: The variation of halo group mass $m_{g}(m_{h})$ for all haloes of the same mass $m_{h}$ at different redshifts $z$ from Illustris-1-Dark simulation. The figure demonstrates the small-scale permanence of the group mass $m_{g}$. Once the statistically steady state is established ($z\leq 4$), the rate of the inverse mass cascade $\varepsilon_{m}$ becomes scale independent. The mass flux into a halo group balances the mass flux out of the same group such that the group mass $m_{g}$ at different redshifts $z$ collapses into a time-independent power law $m_{g}\propto m_{h}^{-\lambda}$ on small mass scales. This is the so-called "propagation range," where the mass cascaded from the smallest (DM particle) scale is propagated to larger scales, that is, a constant group mass $m_{g}$. The cascaded mass is eventually consumed to grow haloes greater than $m_{h}^{*}$ (the "deposition range"). The propagation range gradually extends to large scales ($m_{h}^{*}(a)$ increases over time) due to the continuous inverse mass cascade. The slope $\lambda\approx$1.9 is related to the waiting time, which can be dependent on the initial spectrum.
3.4 The scale-to-scale energy cascade in mass space
In parallel with the mass cascade, there always exists a concurrent flux of energy across haloes of different characteristic scales. This is a unique feature of SG-CFD, as there is only an energy cascade in the incompressible viscous fluid. On cosmological scales, the total energy is not conserved, as implied by Noether’s theorem, because the expanding background lacks time-translation symmetry. As shown in Section B, the cosmic energy monotonically decreases with time in an expanding background. This behavior is analogous to dissipation in fluid turbulence, but here it arises from the expanding background rather than from viscous processes, and it does not entail radiation production. However, effective “dissipation” is not an intrinsic process operating on a global scale. Instead, energy is injected near the scale of the most massive haloes ($m_{h}^{*}$) and subsequently transferred to smaller scales, where it is ultimately “dissipated” by certain small-scale mechanisms. To make this mechanism explicit, we first demonstrate the intrinsic energy loss associated with each halo merging event, which provides a small-scale channel for the “dissipation” of the global cosmic energy budget (Fig. 17). Because merging events continuously remove energy, a sustained energy flux through halo mass space—i.e., an energy cascade—is required to balance the ongoing “dissipation.”
We consider binary merging for two haloes of mass $M_{1}$ and $M_{2}$, and size $r_{h1}$ and $r_{h2}$, respectively. The two haloes are initially well separated. Following merging, the remnant halo has a mass of $M_{3}=M_{1}+M_{2}$ and a size of $r_{h3}$. Assuming that all three haloes share a common bulk velocity $v_{h}$ that is approximately independent of halo mass (Xu:2021-Inverse-and-direct-cascade-of-), the kinetic energy associated with their translational motion is conserved, i.e., $M_{1}v_{h}^{2}+M_{2}v_{h}^{2}=(M_{1}+M_{2})v_{h}^{2}$. We next evaluate the change in the internal (halo) energy due to merging, excluding the large-scale interaction energy between particles belonging to different haloes. By the virial theorem for a potential with an effective exponent $n_{e}$ ($PE\propto r^{n_{e}}$, see Section B), the internal kinetic and potential energies satisfy KE = $n_{e}/2$ PE. Therefore, the change in total internal energy associated with the merging is
$\begin{split}\Delta E_{1}=-\left(1+\frac{n_{e}}{2}\right)\frac{GM_{3}^{2}}{r_{h3}}+\left(1+\frac{n_{e}}{2}\right)\left[\frac{G{M_{1}}^{2}}{r_{h1}}+\frac{G{M_{2}}^{2}}{r_{h2}}\right].\end{split}$ (42)
The density of haloes is related to the mean density $\bar{\rho}_{DM}$ as
where $\Delta_{c}=18\pi^{2}$ is the critical density ratio from the spherical collapse model. Substituting Eq. (43) into Eq. (42), the change in total energy due to the halo merging reads
$\small\begin{split}&\Delta E_{1}=\left(1+\frac{n_{e}}{2}\right)\left(\frac{4}{3}\pi\Delta_{c}\bar{\rho}_{DM}\right)^{\frac{1}{3}}G{M_{1}}^{\frac{5}{3}}\left(1+\gamma^{\frac{5}{3}}-(1+\gamma)^{\frac{5}{3}}\right),\end{split}$ (44)
where $\gamma=M_{2}/M_{1}$ is the mass ratio between two merging haloes. Clearly, the change in total energy due to the merging is always negative or $\Delta E_{1}<0$. Therefore, the structure formation must proceed in a "bottom-up" fashion via halo merging to continuously release system energy, not in a "top-down" fashion that requires an energy penalty. The halo merging provides a viable mechanism to "dissipate" the global-scale cosmic energy. For two merging haloes with a large mass ratio of $\gamma=M_{2}/M_{1}\ll 1$,
$\begin{split}\Delta E_{1}&=-\frac{5}{3}\left(1+\frac{n_{e}}{2}\right)G\left(\frac{4}{3}\pi\Delta_{c}\bar{\rho}_{DM}\right)^{\frac{1}{3}}{M_{1}}^{\frac{2}{3}}M_{2}\\ &=-\frac{5}{3}\left(1+\frac{n_{e}}{2}\right)\frac{GM_{1}}{r_{h1}}M_{2}<0,\end{split}$ (45)
where the change in energy is proportional to $M_{1}^{2/3}M_{2}$. In the smooth mass accretion limit, $M_{2}$ is just the mass of a single merger in Fig. 4. The global-scale energy is continuously "dissipated" by every elementary merging event.
Next, we will consider scale-to-scale energy transfer. In halo mass space, the kinetic energy is cascaded from small to large mass scales at a rate of $\varepsilon_{u}$ (in the same direction as the mass cascade). In contrast, the potential energy cascades in the opposite direction at a rate of $1.4\varepsilon_{u}$ (Fig. 17) to satisfy the virial theorem with $n_{e}=-10/7$ (Section B). Therefore, the total energy is cascaded to small scales at a rate of $0.4\varepsilon_{u}$ and "dissipated" at the same rate. Similarly to the mass flux in Eq. (35), the flux of the specific kinetic energy across haloes reads (also see Eq. (51))
where $\overline{K_{pv}}$ is the mean specific kinetic energy in all haloes greater than $m_{h}$ (defined in Eq. (49)), while $\Lambda_{m}$ is the total mass in all haloes greater than $m_{h}$ (defined in Eq. (37)). For every waiting time of $\tau_{h}$, the energy cascade leads to an increase in the specific kinetic energy of $m_{h}{\overline{K_{pv}}}/{\Lambda_{m}}$ on all scales greater than $m_{h}$.
@@ 3.4 The scale-to-scale energy cascade in mass space
To better describe the energy cascade, we first decompose the kinetic energy of the halo particles into two parts of different nature. In N-body simulations, every halo particle can be characterized by a mass $m_{p}$ and a velocity vector $\boldsymbol{\mathrm{v}}_{\boldsymbol{\mathrm{p}}}$. The particle velocity $\boldsymbol{\mathrm{v}}_{p}$ can be decomposed as (Xu:2023-Maximum-entropy-distributions-of-dark-matter)
@@
namely, the mean velocity of the halo, $\boldsymbol{\mathrm{v}}_{h}=\langle\boldsymbol{\mathrm{v}}_{p}\rangle_{h}$ , and the fluctuation of the velocity, $\boldsymbol{\mathrm{v}}_{p}^{{}^{\prime}}$. Here, $\left\langle\right\rangle_{h}$ represents the average of all particles in the same halo and $\boldsymbol{\mathrm{v}}_{h}$ represents the velocity of that halo. Consequently, the total kinetic energy $K_{p}$ of a given halo particle can be divided into $K_{p}=K_{ph}+K_{pv}$. Here, $K_{ph}=\boldsymbol{\mathrm{v}}_{h}^{2}/2$ (halo kinetic energy) is the contribution from the motion of entire haloes $\boldsymbol{\mathrm{v}}_{h}$ due to the inter-halo interaction of that particle with all other particles in different haloes and all out-of-halo particles. This part of the kinetic energy is related to interactions on large scales in the linear regime.
The other part, $K_{pv}={\boldsymbol{\mathrm{v}}_{p}^{\prime}}^{2}/2$ (virial kinetic energy), is the contribution of the velocity fluctuation $\boldsymbol{\mathrm{v}}_{p}^{\prime}$ due to the intra-halo interaction of that particle with all other particles in the same halo. This part of the kinetic energy is from halo virialization and is due to interactions on a shorter distance and smaller scales in the non-linear regime. Similarly to the energy cascade associated with nonlinear interactions in turbulence, the energy cascade in dark matter flow focused on the cascade of the virial kinetic energy $K_{pv}$ due to nonlinear interactions on small scales.
Similarly to the cumulative mass function $\Lambda_{m}$ in Eq. (37), the cumulative kinetic energies ($\Lambda_{ph}$ and $\Lambda_{pv}$) represent the total kinetic energies $K_{ph}$ and $K_{pv}$ in all haloes greater than $m_{h}$, such that
$\begin{split}&\Lambda_{ph}(m_{h},t)=\int_{m_{h}}^{\infty}M_{h}\left(t\right)f_{M}\left(m,m_{h}^{*}\right)K_{ph}dm,\\ &\Lambda_{pv}(m_{h},t)=\int_{m_{h}}^{\infty}M_{h}\left(t\right)f_{M}\left(m,m_{h}^{*}\right)K_{pv}dm,\\ &K_{ph}=\frac{3}{2}\sigma_{h}^{2}\left(m,t\right)\quad\textrm{and}\quad K_{pv}=\frac{3}{2}\sigma_{v}^{2}\left(m,t\right).\end{split}$ (48)
@@
$\begin{split}\Pi_{pv}(m_{h},t)&=-\frac{\partial}{\partial t}\left(\overline{K_{pv}}\right)=-\frac{\partial}{\partial t}\left(\frac{\Lambda_{pv}}{\Lambda_{m}}\right)\\ &=-\frac{\partial}{\partial t}\int_{m_{h}}^{\infty}\frac{M_{h}\left(t\right)f_{M}\left(m,m_{h}^{*}\right)K_{pv}}{\int_{m_{h}}^{\infty}M_{h}\left(t\right)f_{M}\left(m,m_{h}^{*}\right)dm}dm,\end{split}$ (50)
where $\overline{K_{pv}}$ is defined in Eq. (49), i.e. the specific virial kinetic energy in all haloes greater than $m_{h}$. Similarly to Eq. (36), Eq. (50) describes the rate of transfer of specific virial kinetic energy ($K_{pv}$) from haloes below the scale $m_{h}$ to haloes above the scale $m_{h}$ at a rate of $\Pi_{pv}$.